OpenStudy (anonymous):

the average of five numbers is 9. The average of 7 other numbers is 8. What is the average of all 12 numbers..... These are the options given:

OpenStudy (anonymous):

\[8\frac{ 5 }{ 12 },8\frac{ 3 }{ 4 },8\frac{ 7 }{12 },8\frac{ 11 }{ 12 }\]

hartnn (hartnn):

Since average is 9, sum of 1st 5 numbers=9*5=45 in same way,sum of next 7 numbers=8*7=56 so the sum of 12 numbers=45+56 so average=(45+56)/12 can u simplify this?